differentiate with respect to x : loge3x-3/x
Answer by ptaylor(1894)

differentiate with respect to x : loge3x-3/x
First term:
ln(3x)=ln3+lnx
d/dx(ln3)+d/dx(lnx)=0+1/x=1/x
Second term:
-3/x=-3x^(-1)
d/dx(-3x^(-1))=3x^(-2)
So we have:
1/x+3x^(-2) and we can simplify this:
1/x+3/x^2 multiply each term by (x^2)/(x^2):
(x+3)/x^2
Hope this helps--ptaylor